# Remote electric-field control of antiferromagnetic magnon-polaritons

### v1

@@ Abstract
The control of hybrid light-matter states, specifically magnon-polaritons that emerge from the strong coupling between magnons and cavity photons, remains a key challenge in developing reconfigurable quantum and classical devices. Here, we showcase the ability to remotely control antiferromagnetic magnon-polaritons at room temperature using electric field by integrating a highly birefringent liquid crystal layer into a terahertz Fabry–Pérot cavity containing an antiferromagnetic crystal. Positioned several millimeters from the magnetic material, the liquid crystal allows for adjusting the cavity’s photonic environment through electric field. This adjustment, in turn, influences the coupling strength of a particular cavity mode to the magnon resonance, thereby controlling the extent of magnon dressing by cavity photons. Our approach facilitates dynamic and reversible tuning of magnon-photon hybridization without the need for direct electrical contact or alterations to the magnetic medium. These findings create the conditions for voltage-programmable terahertz magnonic devices and new possibilities for noninvasive control strategies in spin-based information processing technologies.
@@
Magnons are quanta of collective excitation of the electron spins in the magnetically ordered crystals. The ability to transmit information without a charge transfer, thus without heat dissipation, makes them highly attractive to opto-spintronics. Antiferromagnetic magnons have drawn immense attention recently, as they exhibit dynamics in the terahertz (THz) range, much faster than in the case of their ferromagnetic counterparts characterized by the gigahertz (GHz) dynamics. A prerequisite for implementing antiferromagnetic magnons in opto-spintronics is their control by voltage bias, i.e., without electrical current flow. Such type of manipulation of magnons at ambient conditions has been shown exclusively in the case of the multiferroic material BiFeO31, 2 so far. However, the development of a versatile method for voltage-bias control of THz magnons applicable for any type of insulating antiferromagnets at room temperature has been still a challenge.
@@
The strong coupling of light to quasiparticles, such as magnons, excitons, and plasmons, gives rise to new hybrid quasiparticles that can be controlled through their photonic components. In particular, the strong magnon-photon interaction leads to the formation of magnon-polaritons (MPs). These MPs exhibit both light- and matter-like properties, making them highly promising for next-generation high-speed information processing technologies. 3, 4 Over the past decade, cavity magnonics studies have been focused almost entirely on ferromagnetic magnons excited in the GHz range and tuned by the magnetic field.5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15 Only a few reports on modification of the GHz ferromagnetic MPs induced by the electric current16 or voltage bias17, 18 exist. The strong coupling to light of antiferromagnetic magnons in the THz range was demonstrated recently, mostly using temperature tuning.19, 20, 21, 22, 23, 24, 25, 26, 27 The magnetic field and temperature tuning of antiferromagnetic magnon-polaritons is either intrinsically inefficient or too slow from the view of practical applications. Despite being highly desired and crucial for applications in THz opto-spintronics, voltage bias manipulation of THz MPs has not yet been reported.
@@
Liquid crystals provide a highly versatile platform for THz photonic applications, including voltage-induced THz beam manipulation28, 29, phase modulation30 and shifting,31 as well as the tunability of THz metamaterials.32, 33, 34, 35, 36, 37
@@
In this work, we demonstrate the remote electrical control of THz MPs using a liquid crystal cell. We achieve it by integrating, into a Fabry-Pérot cavity, the liquid crystal cell and a 510-m-thick slab nickel oxide (NiO), a room-temperature antiferromagnetic insulator. A separation between the antiferromagnetic layer and the liquid crystal is approximately 2.5 mm. By varying the voltage applied to the liquid crystal, we tune the spatial overlap of the optical modes of the Fabry-Pérot cavity and the antiferromagnet layer (see Figs. Fig. 1(a) and Fig. 1(b)). This allows us to tune the interaction strength between the Fabry-Pérot cavity mode and the magnon above room temperature without external magnetic fields. The presented method of control of magnon-polaritons without direct electrical contact with the magnetic layer can be applied to a wide range of magnetic materials. As such, it is highly attractive for implementation in the THz opto-spintronic systems for control of information transfer.
@@ 1 Results and discussion
We measure the reflection from the Fabry-Pérot cavity composed of NiO and liquid crystal, separated by 2.5 mm (see Fig. 1(a)), using a THz time-domain spectrometer (TDS). We collect the spectra as a function of the NiO temperature () and the bias voltage () applied to the liquid crystal (see the Methods section for details). Two selected time-domain traces for V and V, collected at K, are presented in Fig.1(c). A clear voltage-induced change in the signal is observed around ps, corresponding to a THz pulse reflected from the liquid crystal.
The temperature dependence of reflection spectra of the studied Fabry-Pérot cavity at the voltage bias fixed at V is shown in Fig. 2. Three slightly sloped lines represent the three subsequent optical modes of the Fabry-Pérot cavity (labeled as , , and ). As a magnetic dipole, the magnon line is weakly visible in the spectrum but can manifest through its interactions with the cavity modes. Consistently, as it redshifts with increasing , a series of resonances with the optical modes is observed that enhance the magnon absorption. The strong coupling of the magnon with the middle Fabry-Pérot M2 at around K and THz is the most pronounced, as manifested by the largest anticrossing with the emergence of the lower (MP2) and the upper (MP3) magnon-polariton modes. The splitting between the branches MP2 and the MP3 attains GHz. The interaction of the magnon with the modes M1 and M2 is much weaker. We attribute the strongest coupling observed in the case of the M2 to its largest spatial overlap with the NiO layer. Our calculations of the spatial mode distribution presented in the further text confirm such an interpretation.
To determine MP coupling strength, we fit magnon-polariton energies obtained in a frame of a four-level coupled oscillator model38, 39 to the experimental data. The interaction between the magnon and the three optical modes of the Fabry-Pérot cavity is described by the Hamiltonian:
@@
where Magnon denotes the energy of bare (uncoupled) magnon, Mi (, , ) denote energies of three selected optical modes of the Fabry-Pérot cavity, and (, , ) represent the interaction strength between the magnon and the respective optical mode. We fitted the eigenvalues of to the minima of the -dependent reflection spectra for each applied (see Fig. 2 for the spectra at V). In that way, we obtain the temperature dependencies of the polariton frequencies MPi (, , , ). The frequencies of the uncoupled magnon and Fabry-Pérot cavity modes are determined from the spectra recorded in the conditions, where the magnon-mode coupling is negligible.
@@
To demonstrate remote electric-field control of MP coupling, we apply a voltage bias to the liquid crystal cell. An electric field causes the liquid crystal molecules to reorient. At V, the molecules in the liquid crystal are oriented along the plane of the liquid crystal cell (the configuration known as planar alignment40), as schematically shown in Fig. 1(a) and in the Methods section. When the exceeds a threshold value (), the molecules in the liquid crystal start to reorient continuously with the increasing voltage. At a saturation voltage (), the molecules are aligned perpendicular to the liquid crystal cell plane, as shown in Fig. 1(b). The and depend, in general, on the liquid crystal cell construction; in our case, V and V. The reorientation of the molecules causes a continuous change in the refractive index of the liquid crystal, from the initial value at 0 V bias to at the saturation bias. The change in the refractive index results in modification of frequencies and electromagnetic field spatial distribution of the Fabry-Pérot modes.
@@
The temperature dependence of reflection spectra of the studied Fabry-Pérot cavity for selected values of the varied from 0 V to 30 V are shown in Fig. 3. We focus on the interaction between the magnon and the and modes characterized by the strength and , respectively. With the increasing , we observe an increase of the splitting between polariton branches MP3 and MP4 in the vicinity of the resonance of the uncoupled and magnon transitions occurring at around THz and K. The shift of the splitting center from the resonance results from the contribution from the mode to the wavefunctions of theMP3 and MP4 wavefunctions. The increase of the splitting saturates at V. On the contrary, the increase of the from 0 V to leads to a decrease in the splitting between the branches MP2 and MP3, indicating a reduction of the interaction strength of the magnon and the M2 mode. Frequencies of the uncoupled and modes are only weakly affected by the voltage bias, with the blueshifts not exceeding 5 GHz in the considered voltage range.
@@
In Fig. 4(a), we present reflection spectra in the waterfall format for selected values of at the temperature of 353 K. We clearly see that the interaction of the M3 with the magnon results in the anticrossing and formation of MPs MP3 and MP4 marked with blue and red arrows, respectively. We fitted two Gaussian curves to the MP3 and MP4 transitions to quantify their amplitudes and full widths at half maximum (FWHM). The fits show voltage-induced transfer of amplitude from the MP4 to the MP3 (see Fig. 4(b)). At around 15 V, the resonance between the magnon and occurs, and their interaction is maximum. As expected, at resonance, the amplitudes and the linewidths (see Fig. 4(c)) of MP4 and MP3 are equal.
We determine the coupling strengths by fitting the eigenvalues of to the reflection minima as a function of . Values of and as a function of at K obtained from the fit are presented in Fig. 4(d). Notably, we observed a voltage-induced two-fold increase of and a drop of to around 0.7 of its value at 0 V. Our transfer matrix method calculations of the spatial distribution of the modes (see Fig. 4(e)) indicate that the overlap of the mode increases two times when the voltage is changed from 0 V to 30 V. In the case of the overlap decreases to around 0.65 of the value at 0 V. This confirms that the application of the voltage bias to the liquid crystal for tuning the total spatial overlap of the magnetic field (-field) of the cavity mode with the antiferromagnetic NiO layer provides an efficient means of control of the magnon - mode interaction strength.
@@ 2 Conclusions
We have shown the remote electrical control of THz magnon-polaritons in a Fabry-Pérot cavity embedding an antiferromagnetic NiO layer and a highly birefringent nematic liquid crystal. By applying the voltage to the liquid crystal layer, we tune the liquid crystal refractive index and, in consequence, the spatial distribution of the Fabry-Pérot cavity modes. The change of cavity mode -field overlap with the NiO layer tunes the magnon-photon coupling strength. Importantly, these results have been obtained above room temperature, without external magnetic fields, and with the liquid crystal layer millimeters away from the magnetic material. Reported remote electric-field tunability is a key advancement for remote control of information transfer in the THz opto-spintronics.
@@ 3.1 Experimental setup
The scheme of the experimental setup is presented in Fig. 5(a). We used a NiO single crystal with dimensions of 550.3 mm3, cut in the 111 plane (). The low spin-damping rates of NiO make it an excellent material for MP studies. We tune the magnon frequency by fixing the NiO crystal on a copper mirror using thermo-conductive paste. The mirror is placed on top of a stack of Peltier elements. The NiO temperature is varied in a range of – K, with a step size of K, while the voltage range is – V, with a step size of V. Crystal temperature is monitored with a -type thermocouple sensor placed in the copper mirror. Temperature is stabilized with a software-based PID loop controlling the current supplied to the stack of Peltier elements. The Peltier elements with the mirror and NiO are placed on a kinematic mount and x,y,z stage. The liquid crystal cell is placed on another kinematic mount, allowing precise positioning of the cell to NiO crystal.
We measure THz reflection spectra from the Fabry-Pérot cavity at 0 degree incidence angle. This is achieved using a transceiver antenna integrating source and detector (). We use two parabolic mirrors to focus the incident beam on the Fabry-Pérot cavity and to collect the reflected THz pulses (Fig.5(a)). The THz optical path is filled with dry air to minimize water-vapor absorption. To obtain spectra in the frequency domain, we applied a fast Fourier transform to measured time-domain traces, which are zero-padded from 200 ps to 1000 ps. The absolute value of obtained spectra is normalized to each spectra maximum.
3.2 Liquid Crystal Cell

### v2

# Terahertz magnon-polaritons control using a tunable liquid crystal cavity

@@ Abstract
Strong coupling of light to a collective spin excitation in antiferromagnets gives rise to hybrid modes called magnon-polaritons. They are highly promising for data manipulation and transfer at terahertz rates, much faster than in the case of ferromagnetic magnon-polaritons, which operate at GHz frequencies. Yet, control of terahertz magnon-polaritons by the voltage, i.e. without ohmic dissipation losses, remains challenging. Here, we showcase the ability to remotely control antiferromagnetic magnon-polaritons at room temperature using an electric field by integrating a highly birefringent liquid crystal layer into a terahertz Fabry–Perot cavity containing an antiferromagnetic crystal. Positioned several millimeters from the magnetic material, the liquid crystal allows for electrical manipulation of the cavity’s photonic environment by control of its dielectric constant. This adjustment, in turn, influences the extent of magnon dressing by cavity photons, thereby controlling the vacuum Rabi oscillations of the magnon resonance coupled to a particular cavity mode. Our approach enables reversible tuning of magnon-photon hybridization that can be triggered without direct electrical contact or alteration of the magnetic medium. These findings pave the way for voltage-programmable terahertz magnonic devices and open new avenues for noninvasive control strategies in spin-based information processing technologies.
@@ 1 Introduction
Magnons are collective excitations of the electron spins in the magnetically ordered crystals. The ability to transmit information without a charge transfer, thus without Joule heating, makes them highly attractive to opto-spintronics. Antiferromagnetic, zone-centered () magnons have drawn immense attention recently, as they exhibit dynamics in the terahertz (THz) range, much faster than in the case of their ferromagnetic counterparts, characterized by the gigahertz (GHz) dynamics. A prerequisite for implementing antiferromagnetic magnons in opto-spintronics is their control by voltage bias, i.e., without electrical current flow. Such a type of manipulation of magnons at ambient conditions has been shown exclusively in the case of the multiferroic material BiFeO3PRovillain_NM2010, AKumar_APL2011 so far. However, the development of a versatile method for voltage-bias control of THz magnons applicable for any type of insulating antiferromagnets at room temperature has been still a challenge.
@@
In the strong light-matter coupling regime, vacuum Rabi oscillations, involving the periodic exchange of energy between matter and optical modes, overcome losses, resulting in the emergence of new hybrid modes. The strong coupling of light to quasiparticles, such as magnons, excitons, and plasmons, gives rise to polariton modes that can be controlled through their photonic components. In particular, the strong magnon-photon interaction leads to the formation of magnon-polaritons (MPs). These MPs exhibit both light- and matter-like properties, making them highly promising for next-generation high-speed information processing technologies. KRoux_NC2020, HYuan_PR2022 Over the past decade, cavity magnonics studies have been focused almost entirely on ferromagnetic magnons excited in the GHz range and tuned by the magnetic field.Abe_APL2011, Zhang_PRL2014, Tabuchi_PRL2014, Zhang_NC2015, KGrishunin_ACSPh2018, Everts_PRB2020, Potts_APL2020, Lachance-Quirion_S2020, Li_JAP2020, Bhoi_JAP2021, SGuo_NL2023 Only a few reports on modification of the GHz ferromagnetic MPs induced by the electric currentJTHou_APL2024 or voltage biasKaur:APL2016, Rao:APL2018 exist. The strong coupling to light of antiferromagnetic magnons in the THz range was demonstrated recently, mostly using temperature and magnetic field tuning.KGrishunin_ACSPh2018, PSivarajah_JAP2019, MBialek_PRB2020, MBialek_PRA2021, MBialek_APL2022, ABaydin:PRR2023, MBialek_PRA2023, TGHBlank_APL2023, TElijahKritzell_AOM2023, MBialek_AFM2024, JChen:APL2025 The magnetic field and temperature tuning of antiferromagnetic magnon-polaritons is either intrinsically inefficient or too slow from the view of practical applications. Despite being highly desired and crucial for applications in THz opto-spintronics, voltage bias manipulation of THz MPs has not yet been reported.
@@
Liquid crystals provide a highly versatile platform for THz photonic applications, including voltage-induced THz beam manipulationXFu_ACS2022, ZShen_APLP2024, phase modulationOBuchnev_APL2013 and shifting,LWang_LSA2015 as well as the tunability of THz metamaterials.DShrekenhamer_PRL2013, RKowerdziej_APL2014, RKowerdziej_APL2015, NChikhi_SR2016, RKowerdziej_LC2016, GDENG_OME2021
@@
In this work, we demonstrate the remote electrical control of THz MPs using a liquid crystal cell. We achieve it by integrating, into a Fabry-Perot cavity, the liquid crystal cell and a -m-thick slab of nickel oxide (NiO), a room-temperature antiferromagnetic insulator. A separation between the antiferromagnetic layer and the liquid crystal is approximately 2.5 mm. By varying the voltage applied to the liquid crystal, we tune the spatial overlap of the optical modes of the Fabry-Perot cavity and the antiferromagnet layer (see Figs. Fig. 1(a) and Fig. 1(b)). This allows us to tune the interaction strength between the Fabry-Perot cavity mode and the magnon above room temperature without the need for external magnetic fields. The presented method of control of magnon-polaritons without direct electrical contact with the magnetic layer can be applied to a wide range of magnetic materials. As such, it is highly attractive for implementation in the THz opto-spintronic systems for control of information transfer.
@@ 2 Results and Discussion
We measure the reflection from the Fabry-Perot cavity composed of a NiO slab and a liquid crystal layer, separated by 2.5 mm (see Fig. 1(a)),
@@
using a THz time-domain spectrometer (TDS), as it is shown Fig. 1(c). We collect the spectra as a function of the NiO temperature () and the voltage bias () applied to the liquid crystal (see the Methods section for details). Two selected time-domain traces for V and V, collected at K, are presented in Fig. 1(b), where we mark reflections from consecutive interfaces with numbers 1-5 and dashed lines (see section S1 in the Supporting Information for other raw time-domain traces). A clear voltage-induced change in the signal is observed in the second peak, corresponding to a THz pulse reflected from the liquid crystal cell.
2.1 Temperature Dependence
The temperature dependence of reflection spectra of the studied Fabry-Perot cavity at the voltage bias fixed at V is shown in Fig. 2.
To determine the coupling strength between the magnon and Fabry-Perot modes, we fit magnon-polariton energies obtained in a frame of a four-level coupled oscillator modelWeisbuch:PRL1992, Sciesiek:CM2020, Fas:JPChLett2021 to the experimental data. In general, the coupled oscillator model assumes a coherent exchange of energy between two or more strongly coupled states, resulting in the formation of the two or more polariton branches. The Rabi splitting separating polariton branches is an analogue of a normal-mode splitting observed in the case of, e.g., coupled mechanical oscillators. Here, we assume that the magnon is coupled with three optical modes. We neglect the coupling between the optical modes. This results in the Hamiltonian describing the coupled system as shown below
@@ 2.1 Temperature Dependence
where M denotes the energy of bare (uncoupled) magnon, Pi (, , ) denote energies of three selected optical modes of the Fabry-Perot cavity, and (, , ) represent the interaction strength between the magnon and the respective optical mode. The frequencies of the uncoupled magnon and Fabry-Perot cavity modes are determined from the reflectivity spectra registered at such conditions, where the detuning between the magnon and the cavity modes is much larger than their mutual coupling. The frequencies in the intermediate regions, where the coupling is meaningful, are obtained from interpolation with a polynomial function. We fit the eigenvalues of to the minima of the -dependent reflection spectra for each applied (see Fig. 2 for the spectra at V). In that way, we obtain the temperature dependencies of the polariton frequencies MPi (, , , ). The values obtained from the fit at 0 V are = 2 GHz, = 16 GHz, = 4 GHz.
We assume the strong coupling regime as the condition where the Rabi splitting exceeds the arithmetic average of the linewidths of uncoupled modes. We use Lorentzian to determine the linewidths of the uncoupled cavity modes, , and that of the magnon, .MBialek_AFM2024 With GHz, the – coupling can thus be classified as strong (see section S3 in the Supporting Information) The interactions of the magnon with the modes P1 and do not exhibit resolvable splittings at zero voltage bias. We attribute the strongest coupling observed for the P2 mode to its largest spatial overlap with the NiO layer. Our calculations of the spatial mode distribution, presented later, confirm this interpretation.
To place our strongly coupled system in the context of previous works reporting polariton physics, we compare the obtained coupling strength to the linewidth of the polariton transition. In our case, this ratio amounts to 2.0. The strong coupling between the cavity mode and antiferromagnetic magnon in -Fe2O3 resulted in a ratio of around 6.MBialek_PRA2021 In excitonic systems, based on GaAs, CdTe, or 2D layered semiconductors, the ratio typically remains in the range between 2 and 6.Skolnick1998, Sciesiek:CM2020, Dufferwiel2015 In systems, where plasmonic mode interacts with excitons to form a polariton, the ratio typically remains in the range of up to 5.Torma2014
2.2 Voltage Bias Dependence
@@
To demonstrate remote electric-field control of MP coupling, we apply a voltage bias to the liquid crystal cell. An electric field causes the liquid crystal molecules to reorient. At V, the molecules in the liquid crystal are oriented along the plane of the liquid crystal cell (the configuration known as planar alignmentRKowerdziej_SR2019), as schematically shown in Fig. 1(a) and in the Methods section. When the exceeds a threshold value (), the molecules in the liquid crystal start to reorient continuously with the increasing voltage. At a saturation voltage (), the molecules are aligned perpendicular to the liquid crystal cell plane, as shown in Fig. 1(b). The and depend, in general, on the liquid crystal cell construction; in our case, V and V. The reorientation of the molecules causes a continuous change in the refractive index of the liquid crystal, from the initial value at 0 V bias to at the saturation bias. The change in the refractive index results in modification of frequencies and the electromagnetic field spatial distribution of the Fabry-Perot modes. More details on the voltage bias dependence of uncoupled cavity modes are presented in section S4 of the Supporting Information.
We show in Fig. 3,
the voltage dependence of the reflection spectra of the studied Fabry–Perot cavity at a constant temperature of K. Here, we focus our analysis on the interaction between the magnon and the mode, as its coupling strength shows a pronounced voltage dependence, although the overall coupling to the mode remains considerably stronger. We observe, as a function of voltage, a clear anticrossing of and modes, which is assisted by oscillator strength transfer from to . At around 15 V, the amplitudes of and are equal. This anticrossing occurs when the magnon couples simultaneously to the and modes, with the interaction involving being much stronger. Under these conditions, the weaker – coupling gives rise to a small anticrossing centered at 932 GHz, while the magnon mode (940 GHz) remains strongly detuned from this resonance, as shown in Fig. 3. Similar -dominated interactions are also observed for the and modes in Fig. 2.
To better understand the voltage dependence observed in Fig. 3, we determine the coupling strengths by fitting the eigenvalues of to the reflection minima as a function of for selected values of ranging from 0 V to 30 V, as shown in Fig. 4.
@@
With the increasing , we observe an increase in splittings between polariton branches MP3 and MP4 in the vicinity of the resonance of the uncoupled and magnon transitions occurring at around THz and K. The shift of the splitting center from the resonance results from the contribution from the mode to the MP3 and MP4 wavefunctions. The increase of the splitting saturates at V. On the contrary, the increase of the from 0 V to leads to a decrease in the splitting between the branches MP2 and MP3, indicating a reduction of the interaction strength of the magnon and the P2 mode. We observe that the frequency of uncoupled is almost voltage-independent, while that of mode blueshifts by about 5 GHz with maximum bias. This difference is related to the spatial distribution of each mode in the liquid crystal layer, discussed in detail in Fig. 5(e).
@@
In Fig. 5(a), we present the same reflection spectra as in Fig. 3in the waterfall format for selected values of . We see the anticrossing of MP3 and MP4 marked with blue and red arrows, respectively. We fitted two Gaussians to the MP3 and MP4 transitions to quantify their amplitudes and full widths at half maximum (FWHM). The fits show voltage-induced transfer of amplitude and width from the MP4 to the MP3 (see Fig. 5(b,c)).
Values of and as a function of at K obtained from the fit are presented in Fig. 5(d). Notably, we observed a voltage-induced two-fold increase of and a reduction of to approximately 70% of its value at 0 V. Importantly, these changes originate from large modifications in the spatial distribution of the electromagnetic field within the cavity, as explained in the following paragraphs.
2.3 Simulations
Our transfer matrix method calculations of the mode profiles (see Fig. 5(e) and details in the Methods Simulations section) reveal that the overlap of mode P3 with the antiferromagnetic NiO layer increases by about 50% when the voltage is increased from 0 V to 30 V, while the overlap of P2 decreases slightly from its initial value. This behavior can be attributed to the frequencies of and , which are near the frequency of one of the modes of the free-standing NiO crystal. In this case, altering the refractive index of the liquid crystal layer strongly shifts field amplitude within the NiO layer. This confirms that the voltage bias applied to the liquid crystal enables externally-actuated control of the spatial overlap between the cavity mode magnetic field (-field) and the NiO layer, thereby modulating the magnon-photon coupling strength, without altering the underlying mode frequencies significantly.
The mode remains strongly coupled with the magnon independently of voltage bias. However, the mode, at zero bias, is weakly coupled to the magnon. It enters the strong coupling regime at the voltage bias of around 20 V, when the exceeds the average of the linewidths of the uncoupled magnon and the mode.
In Fig. 6, we simulate reflection spectra with the magnon, reproducing the anticrossings with two cavity modes. This calculation shows qualitatively similar behavior to the experiment.
@@ 3 Conclusions
We show the remote electrical control of THz magnon-polaritons in a Fabry-Perot cavity embedding an antiferromagnetic NiO layer and a highly birefringent nematic liquid crystal. By applying the voltage to the liquid crystal layer, we tune the liquid crystal refractive index and, consequently, the spatial distribution of the Fabry-Perot cavity modes. The change of cavity mode -field overlap with the NiO layer tunes the magnon-photon coupling strength and thus the vacuum Rabi splitting of magnon-polariton modes. Importantly, our results were obtained above room temperature, without external magnetic fields, and with the liquid crystal layer millimeters away from the magnetic material. Reported remote electric-field tunability is a key advancement for remote control of information transfer in the THz opto-spintronics.
@@ 4.1 Experimental Setup
We present the schematics of the experimental setup in Fig. 1(c). We used a NiO single crystal with dimensions of 550.33 mm3, cut in the 111 plane (). The low spin-damping rates of NiO make it an excellent material for MP studies. We tune the magnon frequency by fixing the NiO crystal on a copper plate using thermo-conductive paste, characterized by strong absorption in the THz range. The plate is placed on top of a stack of Peltier elements. The NiO temperature is varied in a range of – K, with a step size of K, while the voltage range is – V, with a step size of V. Crystal temperature is monitored with a -type thermocouple sensor placed in the copper plate. Temperature is stabilized with a software-based PID loop controlling the current supplied to the stack of Peltier elements. The Peltier elements with the plate and NiO are placed on a kinematic mount and an x,y,z stage. The liquid crystal cell is placed on another kinematic mount, allowing precise positioning of the cell relative to the NiO crystal.
We use a commercial TeraFlash proTHz time-domain spectrometer (TOPTICA Photonics). This 80 MHz system employs a pair of photoconductive antennas (emitter and detector) and provides a reliable spectral bandwidth from 0.3 to 2 THz with a frequency resolution of 5 GHz (200 ps scan range) and a signal-to-noise ratio better than 120 dB. We measure THz reflection spectra from the Fabry-Perot cavity at a 0-degree incidence angle. We use three parabolic 1" mirrors to focus the incident beam on the Fabry-Perot cavity and to collect the reflected THz pulses (Fig.1(c)). We use a beam splitter made of Kapton tape to direct the reflected radiation to the detector. We place the beam splitter in the parallel beam formed by the parabolic mirrors. The THz optical path is filled with dry air to minimize water-vapor absorption.
4.2 Fourier analysis
In the time-domain traces, we measured with a photoconductive antenna the electric field of terahertz pulses arriving at the detector. Each trace represents the temporal evolution of the electric field, including the main pulse and subsequent echoes caused by multiple reflections within the sample and between optical components. This direct measurement of the electric field (rather than intensity) preserves both amplitude and phase information, which is essential for accurate spectral analysis. Multiple pulses observed in the time domain correspond to reflections within the sample, and their interference produces characteristic oscillations in the frequency domain. These oscillations appear as deep periodic minima in the amplitude spectrum, which are directly related to Fabry–Perot cavity modes formed between parallel interfaces of the sample. To obtain electric field reflection spectra in the frequency domain presented in the main manuscript, we applied the fast Fourier transform to entire measured time-domain traces (200 ps, zero-padded to 1000 ps) using rectangular windowing, i.e. without applying any additional smoothing or apodization functions, ensuring that all reflections from multiple interfaces are included in the resulting spectra. In the case of Fig. S5 and S6 in the Supporting Information, to filter the frequency domain spectra, we applied rectangular windowing edges set at points where the signal was near zero in time domain. The absolute value of the obtained electric field reflection spectra is scaled to each spectrum’s maximum.
4.3 Liquid Crystal Cavity